### v1

# Graphite: A GPU-Accelerated Mixed-Precision Graph Optimization Framework

Abstract
We present Graphite, a GPU-accelerated non-linear graph optimization framework. It provides a CUDA C++ interface to enable the sharing of code between a real-time application, such as a SLAM system, and its optimization tasks. The framework supports techniques to reduce memory usage, including in-place optimization, support for multiple floating point types and mixed-precision modes, and dynamically computed Jacobians. We evaluate Graphite on well-known bundle adjustment problems and find that it achieves similar performance to MegBA, a solver specialized for bundle adjustment, while maintaining generality and using less memory. We also apply Graphite to global visual-inertial bundle adjustment on maps generated from stereo-inertial SLAM datasets, and observe speed ups of up to $59\times$ compared to a CPU baseline. Our results indicate that our solver enables faster large-scale optimization on both desktop and resource-constrained devices.
I Introduction
Nonlinear optimization is a key component of many estimation problems in computer vision, graphics, and robotics. It is especially prevalent in keyframe-based simultaneous localization and mapping (SLAM) systems, which construct a representation of the environment while also determining a device’s position and rotation. In particular, ORB-SLAM3 [1], a visual-inertial SLAM system, widely employs nonlinear optimization over hypergraphs for a broad array of frontend and backend tasks across its tracking, local mapping, and loop closing threads.
For real-time SLAM, it is necessary to perform these optimizations in a timely manner to remain up to date with changes in the device’s state and surroundings. At the same time, the computational time and amount of memory needed to carry out each optimization may increase as the size of the map grows. This may be especially challenging for resource-constrained platforms when on-device SLAM is just one component of a larger system with several concurrent tasks.
Popular nonlinear optimizers [2, 3, 4] take advantage of multi-core CPU architectures, exploiting the inherent parallelism of sparse linear algebraic operations. However, some platforms used for SLAM [5] are equipped with on-board accelerators such as GPUs, which can be used to perform a variety of learning-based perception tasks efficiently, such as feature extraction, depth-estimation, object detection, and place recognition [6]. GPUs excel at these compute-heavy workloads due to their massively parallel architecture, which also makes them well-suited for nonlinear optimization.
To date, there have been several efforts to speed up existing optimization libraries by identifying and offloading expensive steps to GPUs. More recently, libraries such as Ceres Solver [4] have implemented GPU acceleration for the linear solver step of trust region optimization algorithms. Other work based on g2o [3] focuses on efficiently reducing the size of the linear system to be solved [7] and accelerating numerical differentiation for computing Jacobian matrices [8]. Yet, while it is possible to improve the performance of existing libraries by offloading individual steps to the GPU, it introduces additional overhead from allocating GPU memory and transferring data between steps.
An alternative approach is to perform the entire optimization on the GPU. Some methods use Python [9] or a domain-specific language (DSL) [10, 11] to allow users to write simple descriptions of nonlinear optimization problems which are then used to automatically generate efficient solvers using complex code transformations. Others leverage machine learning frameworks to support batched optimization [12] and differentiable optimization [13, 14] for training learning-based tasks such as feature extraction and feature matching. In addition, many solvers are specialized for specific classes of problems such as bundle adjustment [15], achieving large speedups.
However, existing GPU-accelerated solvers have several limitations which make them challenging to adapt for applications such as SLAM. They often model optimizable variables as simple numeric data types (e.g. a float, double, vector of doubles, etc.) or only provide data types for a specific optimization problem (e.g. a class representing an SE(3) transformation). Meanwhile, SLAM systems may represent variables as complex classes which consist of other classes. For example, an optimizable pose class may consist of multiple SE(3) transformations and cameras [1]. Existing solvers cannot represent these classes with a vector of numbers or predetermined types because they cannot model their complex behaviours and data dependencies. Additionally, researchers develop real-time SLAM systems in a specific language of their choice, e.g., C++, making it difficult to use DSLs or Python-based optimizers, since they require optimizable data types and functionality to be faithfully reimplemented for another language. A third challenge is that GPU memory is often limited and shared between multiple tasks such as image processing and model inference. Moreover, existing libraries may demand additional GPU memory for operations on sparse matrices and linear solver workspace allocations [16]. Even worse, embedded platforms often have no dedicated GPU memory, so the CPU and GPU must compete for the same memory.
In this paper, we present Graphite, a GPU-accelerated nonlinear optimization framework for estimation problems, which implements several techniques for balancing runtime performance and GPU memory usage, to enable large-scale optimization for desktop and embedded scenarios. It allows for mixed-precision solving using 64-bit, 32-bit, and 16-bit floating-point precisions, enabling faster and more memory-efficient optimization of graphs. To further reduce memory usage, the library supports dynamically computed Jacobians for matrix-free methods, as well as an automatic differentiation method with equivalent memory overhead to analytic differentiation. In addition, the iterative linear solver is aware of the structure of the graph, bypassing explicit sparse matrix formats and their associated memory costs. To support real-time SLAM and odometry, the framework provides a CUDA C++ interface, and uses a batching model which supports in-place optimization, meaning that users can use their own data types, without having to transform them into a solver-compatible format. This avoids memory usage and runtime overhead from unnecessary copying, as well as explicit data transfers on platforms which share CPU and GPU memory.
To demonstrate the effectiveness and flexibility of our framework, we evaluate our solver on bundle adjustment problems, and find it performs comparably to MegBA [15] while using up to 78% less GPU memory. We also reimplement visual-inertial bundle adjustment on the GPU inside ORB-SLAM3, which can take hundreds of seconds on the CPU for sufficiently large problems, and demonstrate a speed up of up to $59\times$ on maps generated by processing stereo-inertial SLAM datasets [17], while using under 1 GiB of GPU memory on the desktop. The code is available at https://github.com/sfu-rsl/graphite.
To summarize our contributions:
  • We design and implement a general mixed-precision optimization framework for on-device estimation problems which performs all major optimization steps on a GPU.
  • We apply our framework to global visual-inertial bundle adjustment inside a SLAM system, which consists of 7 types of constraints and 5 types of variables.
  • We evaluate our solver across different precisions and differentiation modes, using well-known bundle adjustment and SLAM datasets.
II Background
In this section, we provide a brief discussion on nonlinear optimization over constraints, the Levenberg-Marquardt algorithm implemented by our framework, visual-inertial optimization, and relevant GPU programming concepts.
Optimizing Over Constraints: Some estimation problems can be modelled as an optimization over constraints. Each constraint represents a relationship between estimated parameters, an observation, and the error. The goal is then to find the parameters which minimize the total error across all constraints ($\mathcal{C}$). This may be formulated as
$\footnotesize\text{arg}\min\limits_{x}\frac{1}{2}\sum_{\langle i,j\rangle\in\mathcal{C}}r_{ij}^{\top}\Sigma_{ij}^{-1}r_{ij}$ (1)
where $x$ is the parameter vector, $r_{ij}$ is the residual at $x$ associated with a constraint between variables $i$ and $j$ and $\Sigma_{ij}$ is the covariance matrix for the constraint. Alternatively, the quadratic form can be written as
$\footnotesize\text{arg}\min\limits_{x}\frac{1}{2}r^{\top}\Sigma^{-1}r.$ (2)
These problems may be expressed using an optimization framework [2, 3, 4] and solved using algorithms such as Powell’s dog leg or Levenberg-Marquardt.
Levenberg-Marquardt for Nonlinear Least Squares: Given the non-linear optimization problem in Eq. 1, the Levenberg-Marquardt algorithm (LM) can be used to iteratively find new parameters to drive the error towards a local minimum. LM combines aspects of Gauss-Newton and gradient descent to achieve better convergence with the use of a damping factor. It exploits the first-order Taylor expansion
$\footnotesize f(x+\Delta x)\approx f(x)+\frac{\partial f(x)}{\partial x}\Delta x.$ (3)
Each iteration then solves
[…]
$\footnotesize(J^{\top}\Sigma^{-1}J+\lambda I)\Delta x=-J^{\top}\Sigma^{-1}r.$ (5)
This can be more simply rewritten in terms of the regularized Hessian approximation $H$ and gradient $b$ as
$H\Delta x=-b.$ (6)
Solving this linear system in each iteration yields $\Delta x$, which represents a candidate parameter step to reduce the total error. Here, $H$ is a sparse positive semi-definite block matrix.
Visual-Inertial Optimization: Given a set of keyframes $\kappa$ and $l$ map points, the goal of visual bundle adjustment is to find new parameters which better satisfy camera reprojection error constraints. The cost may be formulated as Eq. 7  [1]
$\footnotesize\sum_{j=0}^{l-1}\sum_{i\in\kappa^{j}}\rho_{\text{Huber}}(\|r_{ij}\|_{\Sigma_{ij}^{-1}}^{2})$ (7)
where $r_{ij}$ is the residual vector of the reprojection constraint between keyframe $i$ and observed map point $j$. In other words, $r_{ij}$ is the difference, in 2D coordinates, between the measured observation of a map point in a keyframe image and its modelled observation computed from the camera’s projection function using current keyframe and map point parameters. To reduce sensitivity to feature mismatches between keyframes and potential outliers, a robust Huber kernel $\rho_{\text{Huber}}$ is used [1]. Visual-inertial bundle adjustment introduces additional terms to the optimization for inertial parameters. For k+1 keyframes, this is represented as [1]
$\footnotesize\sum_{i=1}^{k}\|r_{I_{i-1,i}}\|_{\Sigma_{I_{i,i+1}}^{-1}}^{2}+\sum_{i=1}^{k}\|r_{b_{i-1,i}}\|_{\Sigma_{b_{i,i+1}}^{-1}}^{2}$ (8)
where $r_{I_{i-1,i}}$ is the inertial residual between consecutive keyframes, representing the discrepancy between preintegrated IMU measurements and the change computed from keyframe parameters. These parameters include rotation, velocity, and position. Similarly, $r_{b_{i-1,i}}$ is the accelerometer and gyroscope bias change between consecutive keyframes.
GPU Programming: Modern GPUs are based on massively parallel Single-Instruction, Multiple-Thread (SIMT) architecture, where threads execute common instructions in groups. CUDA [18] is a programming model which enables writing specialized functions, called kernels, which can be dispatched onto a GPU in order to processes a batch of similar tasks in parallel. CUDA uses a hierarchical execution model where threads are grouped into 3D thread blocks. Thread blocks are further organized into a grid. When a kernel is launched, threads and blocks are assigned identifiers which can be queried by a task to decide inputs, outputs, and behaviour. Threads can then communicate with each other using shared memory and built-in functions. Additionally, multiple kernels can be launched in different streams, where each stream is a sequence of commands, to allow their executions to overlap on the GPU.
III Related Work
As our work proposes a GPU-based approach for non-linear least squares optimization for problems such as bundle adjustment, we review literature on these topics.
Although there are several versatile libraries for least squares optimization in robotics and computer vision [2, 3, 4, 19, 20, 21], they are primarily CPU-based and they generally limit GPU acceleration to the linear solver stage [4]. Our framework differs by performing the entire optimization on the GPU, to enable greater parallelization for variable and constraint operations, and avoid data transfers between intermediate steps to improve performance.
There are also a number of GPU-accelerated solvers based on domain-specific languages (DSLs) [9, 10, 11] or Python-based machine learning frameworks [12, 13, 14]. However, these approaches may result in code duplication, lack the expressiveness needed to model complex types, or involve explicit language interoperation, depending on the application. Our framework avoids this by allowing problems to be described directly using user-defined types, preventing code duplication, and enabling easier prototyping.
Several prior works also explore specialized implementations for bundle adjustment on the GPU, applying direct or iterative solving techniques [22, 23, 24, 25, 15, 26, 27]. In practice, these libraries are designed and implemented for a specific type of visual bundle adjustment which only uses camera poses, 3D points, and binary constraints. Meanwhile, SLAM and odometry systems often add several new constraints by incorporating inertial information into the optimization [1, 28], where some constraints have as many as six variables each. At the same time, they may perform other optimization tasks such as motion-only bundle adjustment and pose graph optimization [1, 28]. Our framework resolves this limitation by supporting user-defined types along with unary, binary, and n-ary GPU constraints.
While prior GPU-accelerated solvers utilize code generation and leverage existing machine learning frameworks, we adopt a design inspired by libraries such as g2o and Ceres Solver, to support the level of functionality needed for visual-inertial SLAM [1] and odometry systems [28], discussed further in Section V. The CUDA C++ interface enables the ability to leverage existing data structures within these systems for optimization. This approach also enables in-place optimization on platforms with full unified memory support, so that variables need not be explicitly transferred. Additionally, we support the ability to mix higher and lower floating-point precisions to reduce GPU memory usage.
IV Design
Our optimization framework, based on the Levenberg-Marquardt algorithm (Section II), allows for an optimizable graph to be constructed from user-defined types and unary, binary, and n-ary constraints, to support a wide variety of optimization problems. Its design is guided by the need to minimize the amount of time needed for optimization while also using as little GPU memory as possible, which is crucial for resource-constrained systems. Next, we describe aspects of our framework that are relevant to this approach.
IV-A Overview
Graphite implements graph-based optimization, where all major steps run on the GPU. It allows optimizable variables (vertices) to be represented by classes defined by the user, which avoids the need to first transform variables into a specific format for optimization, a step which is often required by other libraries. These variables can then be iteratively refined in-place on the GPU, using a batching method described in Section IV-B. As is standard with other frameworks, Graphite supports fast analytic differentiation (Section IV-C) and convenient automatic differentiation (Section IV-D) for computing Jacobian matrices used by the optimization algorithm. To perform the optimization process quickly and in a memory-efficient manner, Graphite also supports mixed-precision solving (Section IV-E).
IV-B Batching
Suppose a user wants to optimize noisy 2D points so that they lie on the circumference of a circle. A typical approach to offloading work onto a GPU is to organize similar tasks into batches, and process tasks within a batch concurrently across numerous threads in order to take advantage of the massively parallel architecture. However, existing graph-based optimizers store different types of vertices and constraints indiscriminately together in the graph. This is not ideal for batching work on the GPU, because each thread processing an item (a vertex or constraint) may execute a different code path, which is detrimental to performance [29]. Therefore, to enable a GPU-friendly approach for graph-based optimization, Graphite allows users to construct an optimizable graph from collections of vertices and constraints of the same type. This enables GPU threads to process batches of identical items, which is better for performance [29]. We represent these collections using descriptors, which we describe next.
We illustrate how the user models the point optimization problem in Graphite, using mixed-precision where the user selects single precision for the descriptors, and 16-bit brain floating point for the linear system. First, the user creates a vertex descriptor for a collection of noisy points for the optimization problem, as shown in Figure 1. They specify single precision for the descriptors (Line 1) and 16-bit brain floating point for the linear system (Line 2). In Lines 4-6, they create a descriptor, reserve memory for the desired number of vertices, and add the descriptor to the graph. Then from Lines 8-10, they add a pointer for each vertex to the descriptor, and associate it with an identifier. In creating this descriptor, the user defines a batch of vertices which is processed efficiently on the GPU across many threads.
using FP = float; // vertex, constraint precision
using SP = __nv_bfloat16; // linear system precision
[…]
point_desc.reserve(num_vertices); // Reserve memory for num_vertices
graph.add_vertex_descriptor(&point_desc); // Add to graph
[…]
point_desc.add_vertex(vertex_id, &points[vertex_id]); // Add points
Figure 1: Creating a vertex descriptor for 2D points.
Next, they create a descriptor for the constraints in Figure 2. As in the previous example, they create a descriptor, reserve memory for the desired number of constraints (one constraint for each vertex), and add the descriptor to the graph (Lines 1-3). They initialize the default loss function on Line 4, and then create the desired number of constraints (Lines 6-8). When creating each constraint, the user specifies the identifiers of the vertices involved, the observation (in this case, the radius of the circle), the precision matrix (nullptr defaults to the identity matrix), the data for non-optimizable variables (not used), and the loss function parameters. Again, by constructing this descriptor, the user enables each constraint to be processed concurrently and efficiently on the GPU.
auto factor_desc = CircleFactor<FP, SP>(&point_desc); // Create descriptor
factor_desc.reserve(num_vertices); // Reserve memory for constraints
graph.add_factor_descriptor(&factor_desc); // Add descriptor to graph
const auto lf = DefaultLoss<FP, 1>(); // loss function
[…]
factor_desc.add_factor({vertex_id}, radius, nullptr, 0, lf);
Figure 2: Creating a descriptor for constraints.
At compile time, the properties of each descriptor are used to specialize each CUDA kernel for operations over a specific constraint or vertex type, so that each thread carries out the same type of computation, which improves performance. For a vertex descriptor, these properties include information such as the underlying data type (class), the size of its parameter block, a method to get its parameter block, and a method to update each vertex for a new step. Figure 3 shows how the user defines the properties for the 2D point descriptor. They specify the underyling vertex data type as a $2\times 1$ Eigen [30] matrix (Line 2, Line 7), the length of the parameter block as 2 (Line 6), the method to get the parameter block (Lines 9 to 11), and the update method (Lines 13-16).
// Point definition
template <typename T> using Point = Eigen::Matrix<T, 2, 1>;
[…]
template <typename T, typename S> struct PointTraits {
static constexpr size_t dimension = 2;
using Vertex = Point<T>;
[…]
return vector_to_array<T, dimension>(vertex);
[…]
Eigen::Map<const Eigen::Matrix<T, dimension, 1>> d(delta);
vertex += d;
[…]
template <typename T, typename S>
using PointDescriptor = VertexDescriptor<T, S, PointTraits<T, S>>;
Figure 3: Properties of a vertex descriptor for a 2D point.
In general, properties for a constraint descriptor include which vertex descriptors are involved, the data types (classes) of the observations and any non-optimizable data, the loss function used, and the differentiation mode (discussed in Section IV-C and Section IV-D). These properties also specify how to compute the residual for a constraint, and how to compute the Jacobians when using analytic differentiation (Section IV-C). Figure 4 shows this for the point optimization problem. The user specifies a residual of size 1 (Line 3), specifies the point vertex descriptor (Line 4), chooses the default loss function (Line 7), and chooses to use automatic differentiation (Line 8), so no Jacobian evaluation function is specified. Additionally, Lines 10-18 tell the descriptor how to compute the residual based on the squared distance between the point and the radius.
We specify properties of descriptors in this way because it lets users declare the behaviour of their types and automatically generate optimized GPU code, without relying on runtime lookups (as observed in some CPU-based optimizers [3]) to decide how each item in the graph is processed.
// Factor traits for the circle constraint
template <typename T, typename S> struct CircleFactorTraits {
static constexpr size_t dimension = 1;
using VertexDescriptors = std::tuple<PointDescriptor<T, S>>;
using Observation = T;
using Data = unsigned char;
using Loss = DefaultLoss<T, dimension>;
using Differentiation = DifferentiationMode::Auto;
[…]
hd_fn static void error(const D *point, const T *obs, D *error,
const std::tuple<Point<T> *> &vertices,
const Data *data) {
auto x = point[0];
auto y = point[1];
auto r = obs[0];
error[0] = x * x + y * y - r * r;
[…]
using CircleFactor = FactorDescriptor<T, S, CircleFactorTraits<T, S>>;
Figure 4: Properties of a constraint checking a point’s distance from the circumference of a circle with radius $r$.
After the user finishes adding descriptors to the graph and populating them with information about each vertex and constraint, they pass the graph to an optimization algorithm, along with a linear solver (discussed further in Section IV-E). The user optimizes the graph for the point optimization example in Figure 5. First, they decide to fix the value of the last point they created (Line 1) and decide to disable the constraint which was created for the third point by changing its level to 1, as in g2o [3] (Line 2). Next, the user configures the iterative linear solver, by selecting an identity matrix for the preconditioner (Line 3) and then instantiating the solver (Line 4) with their desired parameters. They choose 10 iterations for Levenberg-Marquardt (Line 6), and create one stream (Line 7), since each constraint only has one vertex. On Line 8, they set the optimization level to 0, meaning that any constraints with a level higher than 0, such as the constraint on Line 2, will be excluded from the optimization. Finally, they perform the optimization by calling Levenberg-Marquardt in Lines 9-10.
point_desc.set_fixed(num_points - 1, true); // Set the last vertex as fixed
factor_desc.set_active(2, 0x1); // Disable third constraint for point 2
IdentityPreconditioner<FP, SP> preconditioner;
PCGSolver<FP, SP> solver(50, 1e-6, 10.0, &preconditioner);
[…]
StreamPool streams(1); // chose 1 stream for 1 vertex per constraint
constexpr uint8_t optimization_level = 0;
optimizer::levenberg_marquardt<FP, SP>(&graph, &solver, iterations, 1e-6,
optimization_level, streams);
Figure 5: Example of optimizing noisy 2D points.
IV-C Analytic Differentiation
In graph-based optimization, constraints may involve one or more vertices, each of which generate a Jacobian matrix. A typical way to compute Jacobians quickly is by evaluating an analytic (closed-form) expression. Graphite allows Jacobians to be evaluated directly using analytic expressions defined in a user-written function. These expressions are derived by hand or generated using a symbolic toolkit [21, 31].
A typical approach on the CPU is to compute each matrix for a constraint sequentially [3, 4]. However, to take advantage of the many GPU threads available, Graphite computes each Jacobian concurrently in separate streams (we describe streams in Section II). Additionally, Graphite allows these Jacobians to be evaluated dynamically [22], at the point of computation (e.g. matrix-vector multiplication), to enable matrix-free solving methods which use less GPU memory.
IV-D Automatic Differentiation
Another method to compute Jacobians is through forward mode automatic differentiation. For example, Ceres Solver computes the gradient of a residual function using dual numbers, which consist of a real component $a$, for the computed value, and an infinitesimal component $b\varepsilon$, where $\varepsilon^{2}=0$, for the corresponding partial derivative [4]. To extend this for multiple inputs, Ceres implements a data structure called a Jet, which replaces the infinitesimal component with a vector [4]. Ren et. al [15] extend this concept to the GPU by proposing JetVectors, which adopt a structure of arrays data layout to enable coalesced memory transactions for better performance. However, this design is intended for optimization across multiple devices and may require more GPU memory than available on a single device [15].
To handle environments with less available memory, we avoid allocating storage for dual numbers entirely. Instead of using Jets or JetVectors, when computing the residual of a constraint, each kernel replaces the input parameters with vectors of dual numbers, one for each vertex. Each CUDA thread then calculates the column of a Jacobian matrix (Figure 6), by computing the residual using these vectors. The column to be computed is changed by setting one partial to 1, and the others to 0. As shown in Section V, this uses the same amount of memory as analytic differentiation, allowing Jacobians to be computed efficiently and conveniently without first needing to derive an expression.
Figure 6: A CUDA thread calculates $\frac{\partial f}{\partial x_{3}}(\vec{x})$ by replacing the input with a vector of dual numbers and writes it to the corresponding Jacobian column.
IV-E Mixed-Precision Solver
To enable mixed-precision solving, which enables the optimization to use less memory, Graphite supports the customization of floating point precisions for descriptors and the linear system (Eq. 6). Essentially, these two precisions can be customized according to trade-offs between convergence, runtime performance, and memory usage. A typical approach would be to use a higher precision (double or single precision) for the descriptors to maintain numerical stability, and an equal or lower precision (single precision or brain floating point), to reduce the amount of memory needed to store the Jacobian matrices [32]. In contrast, existing solvers support only one precision at a time [15], or only allow casting the linear system to single precision [4]. As shown in Section V, using lower precisions greatly reduces the overall optimization time and amount of GPU memory required.
As described in Section II, each iteration of Levenberg-Marquardt optimization computes a parameter step $\Delta x$. To compute this quickly on the GPU, we apply preconditioned conjugate gradients (PCG) on the Hessian, with implicit evaluation using the Jacobian matrices [22], since this does not require storing the Hessian, and comprises of algebraic operations which are largely amenable to parallelization. However, directly using PCG with lower precision types can result in poor convergence. To mitigate this, we adopt techniques which are known to help with solving larger systems. These include, clamping values along the Hessian diagonal [4], rescaling the column of each Jacobian block by the square-root inverse of the corresponding Hessian diagonal value [4, 32], and normalizing the PCG residual in each iteration [33].
In applying PCG, to improve convergence, we also use a block Jacobi preconditioner [34], because it is relatively simple to compute and does not require much memory due to its sparse block-diagonal matrix structure. We have observed that storing this preconditioner at too low of a precision can also result in convergence issues. Thus, when a 16-bit precision type such as bfloat16 is specified for the linear system, we compute and store this preconditioner at single or double precision, according to the precision of the graph.
Additionally, since the PCG solver does not require storing the Hessian matrix, we avoid explicitly constructing sparse matrices for the optimization problem. Instead, when performing an operation such as matrix-vector multiplication, the kernel uses the structure of the graph itself. This eliminates additional memory and runtime costs associated with converting and storing matrices in different formats, although each kernel must now be aware of the graph structure and whether a variable is fixed or constraint is inactive.
V Evaluation
Machine Specs
12-core Intel Core i7-12700K CPU @ 4.9 GHz
Desktop 10496-core NVIDIA RTX 3090 GPU @ 1.9 GHz
64 GB RAM
6-core ARM Cortex-A78AE CPU @ 1.7 GHz
Orin Nano 1024-core NVIDIA Ampere GPU @ 1.1 GHz
8 GB RAM
TABLE I: Evaluation machine specifications.
Problem Metric FP64 FP32 FP32-BF16 Ceres-CGNR MegBA-FP64 MegBA-FP32
MSE 0.43 0.43 0.43 0.43 0.43 0.43
Dubrovnik-16 Time 0.22 0.11 0.12 1.64 0.17 0.16
GPU Memory 298 282 278 336 406 352
MSE 1.67 1.67 1.68 1.67 1.67 1.67
Trafalgar-21 Time 0.11 0.069 0.068 0.82 0.098 0.088
GPU Memory 280 274 270 318 326 298
MSE 0.85 0.85 0.85 0.84 0.84 0.84
Ladybug-49 Time 0.063 0.041 0.046 0.66 0.086 0.074
GPU Memory 278 274 270 316 326 300
MSE 0.67 0.67 0.67 0.67 0.66 0.67
Venice-1778 Time 6.00 2.80 2.87 43.71 6.96 6.15
GPU Memory 1944 1190 950 2160 7810 4528
MSE 1.13 1.13 1.14 1.14 1.12 1.13
Final-4585 Time 9.62 4.69 3.75 73.49 13.19 9.82
GPU Memory 3102 1822 1386 3546 13882 7960
MSE 1.50 1.50 1.50 1.51 OOM OOM
Final-13682 Time 26.60 11.65 11.67 210.64 OOM OOM
GPU Memory 9298 5212 3828 10640 OOM OOM
TABLE II: Time (seconds), mean squared error, and GPU memory usage (MiB) for BAL experiments using different precisions (desktop).
Problem Metric Analytic Dynamic Auto
MSE 1.50 1.50 1.50
Final-13682 Time 26.57 417 38.81
(FP64) GPU Memory 9298 3988 9298
MSE 1.50 1.50 1.50
Final-13682 Time 10.97 25.22 11.46
(FP32) GPU Memory 5212 2556 5212
MSE 1.50 1.50 1.50
Final-13682 Time 10.64 24.48 11.27
(FP32-BF16) GPU Memory 3828 2500 3828
TABLE III: Time (seconds), mean squared error, and GPU memory usage (MiB) for across differentiation modes (desktop).
Sequence Initial $\chi^{2}$ Graph Size Metric g2o-LDLT g2o-PCG FP64 FP32 FP32-BF16
3131 KF Final $\chi^{2}$ $3.0918\times 10^{5}$ $4.4965\times 10^{5}$ $4.3850\times 10^{5}$ $4.6842\times 10^{5}$ $5.9817\times 10^{5}$
outdoors1 $1.0803\times 10^{6}$ 58886 MP Time 1424.40 224.08 4.96 1.71 1.18
(Desktop) 348548 Co GPU Mem. 678 580 542
2032 KF Final $\chi^{2}$ $2.9105\times 10^{5}$ $3.5100\times 10^{5}$ $3.5012\times 10^{5}$ $3.4090\times 10^{5}$ $4.8966\times 10^{5}$
outdoors2 $7.3020\times 10^{5}$ 57641 MP Time 698.68 147.57 2.80 3.56 1.10
(Desktop) 338364 Co GPU Mem. 656 564 530
1595 KF Final $\chi^{2}$ $2.5617\times 10^{5}$ $3.0916\times 10^{5}$ $3.0354\times 10^{5}$ $0684\times 10^{5} 3.3$$3927\times 10^{5}$
outdoors3 $6.4905\times 10^{5}$ 36536 MP Time 2957.40 99.69 1.69 1.32 1.13
(Desktop) 272307 Co GPU Mem. 636 562 534
2483 KF Final $\chi^{2}$ $2.2494\times 10^{5}$ $2.9363\times 10^{5}$ $2.9112\times 10^{5}$ $2.9348\times 10^{5}$ $3.6397\times 10^{5}$
outdoors1 $7.1589\times 10^{5}$ 48621 MP Time 2411.40 512.58 122.61 93.10 27.90
(Orin Nano) 278661 Co GPU Mem. 5673, 306 5462, 95 5433, 66
1972 KF Final $\chi^{2}$ $2.1432\times 10^{5}$ $3585\times 10^{5} 3.3$$3731\times 10^{5}$ $3.5705\times 10^{5}$ $3.5889\times 10^{5}$
outdoors2 $7.3551\times 10^{5}$ 33831 MP Time 1141.40 345.46 53.63 32.58 29.73
(Orin Nano) 213125 Co GPU Mem. 5128, 328 5031, 231 5017, 217
1589 KF Final $\chi^{2}$ $1.6574\times 10^{5}$ $1.9508\times 10^{5}$ $1.9439\times 10^{5}$ $2.0736\times 10^{5}$ $2.1015\times 10^{5}$
outdoors3 $5.3520\times 10^{5}$ 24524 MP Time 2166.1 355.34 93.71 33.53 29.64
(Orin Nano) 174070 Co GPU Mem. 4674, 282 4473, 81 4449, 57
TABLE IV: Results of full-inertial bundle adjustment on generated maps. Time given in seconds and GPU memory in MiB.
We evaluate Graphite in terms of the error reduction, runtime performance, and GPU memory usage. First, we compare the solver in different precision modes, using problems from Bundle Adjustment in the Large (BAL) [34], a popular dataset for least squares and bundle adjustment solvers. Next, we compare analytic, dynamic, and automatic differentiation modes. Lastly, we run full-inertial bundle adjustment with Graphite inside ORB-SLAM3 using maps generated from outdoors sequences in TUM-VI, a well-known dataset for evaluating visual-inertial SLAM systems. We run these experiments using a desktop machine equipped with a discrete graphics card, and an NVIDIA Jetson Orin Nano Developer Kit, an embedded robotics platform, to assess the performance of Graphite in high-performance and resource-constrained environments (specifications in Table I).
V-A Mixed-Precision Modes
We evaluate Graphite on BAL datasets using double precision (FP64), single precision (FP32), and single precision mixed with bfloat16 (FP32-BF16) on the desktop machine. We use bfloat16 rather than 16-bit half precision, because it can approximately represent the same range as single precision. For comparison, we also include two GPU-accelerated solvers, Ceres Solver [4], a versatile solver commonly used for SLAM, configured for CUDA CGNR (conjugate gradients on the normal equations), which like our solver, also uses implicit Hessian evaluation, and MegBA [15], a high-performance solver for bundle adjustment with support for single and double precision solving. Where available, we use analytic Jacobians, which have the least runtime overhead. We generate analytic Jacobians for our solver using wrenfold [31] for symbolic differentiation. To allow each optimizer sufficient opportunity to converge, we use a maximum of 50 iterations. For PCG in MegBA, the maximum number of iterations is set to 100, while 10 is used for the other configurations, since they solve the full system rather than a reduced one, which involves more work per iteration. As shown in Table II, the solver is able to use substantially less memory when using BF16 to store the Jacobian matrices, at the cost of slightly worse convergence. For runtime, FP64 takes the longest, while FP32 and FP32-BF16 take a comparable amount of time. Compared to Ceres-CGNR, Graphite takes less time since it accelerates constraint-specific calculations as well. Although MegBA achieves slightly lower MSE in a few instances, Graphite uses $\frac{1}{4}$ the memory for large problems, since it avoids storing Hessian matrices, and receives a greater speedup when using single precision. These results highlight the runtime and memory advantages of performing the entire optimization on the GPU and directly utilizing the graph structure to avoid constructing intermediate matrices.
V-B Differentiation Modes
The performance of our solver on the desktop using analytic, dynamic, and automatic differentiation modes is given in Table III for Final-13682, the largest dataset in BAL. Automatic differentiation allocates the same amount of memory as analytic differentiation, taking marginally more time at lower precisions, and significantly more at FP64. Meanwhile, the dynamic mode requires the least amount of GPU memory since Jacobians are not stored, although precision matrices are still used. While the dynamic mode exhibits a larger slowdown for FP64, this can be mitigated by switching to lower precisions. Our results show that automatic differentiation can be achieved with relatively low memory and runtime overhead for lower precisions, enabling rapid prototyping of optimizable constraints, while dynamically computed Jacobians enable more memory efficient solving at the cost of higher runtimes.
V-C Case Study: Visual-Inertial Bundle Adjustment
We apply Graphite to full-inertial bundle adjustment in ORB-SLAM3, which introduces several additional variables and constraints to visual bundle adjustment (Section II).
Implementation: To use Graphite for visual-inertial bundle adjustment, it is necessary to reimplement the vertices and constraints involved in the optimization. The original data types in ORB-SLAM3 cannot be used as-is, since some contain extraneous data, involve dynamic memory allocations which must be manually transferred from the CPU to the GPU, or implement methods which are not suitable for GPU execution (e.g. a method which accesses a resource shared between different CPU threads). Therefore, we implement lightweight template data types as a replacement, which contain only the subset of data and functionality needed for visual-inertial optimization. These include data types for the combined IMU camera pose vertex, the cameras (pinhole and fisheye), and the preintegrated IMU measurements. The remaining vertex data types for the map points, velocities, gyro biases, and accelerometer biases are $3\times 1$ vectors, so we directly represent these using Eigen [30]. To use these data types with Graphite, we create five vertex descriptors.
To model the optimization constraints, we create eight descriptors for monocular and stereo reprojection errors, the inertial residual (with and without Huber loss), gyro and accelerometer bias residuals, and bias priors. The residual and analytic Jacobian expressions for each constraint are mostly a direct translation of the original ORB-SLAM3 implementation, although we found that some functionality does not behave as expected on the GPU. For example, ORB-SLAM3 normalizes a rotation matrix $R$ by first computing its SVD decomposition and then computing $R=UV^{\intercal}$, but this fails when done in a kernel, so we replace it with a Newton-Schulz iteration. For Lie groups, we use Sophus [35], as in the original implementation.
When constructing the graph, to avoid overhead from reallocating memory, we predetermine the size of the optimization, reserve memory for each descriptor, and preallocate unified memory to store the optimizable variables.
Experiment: Next, large visual-inertial maps are generated by processing TUM-VI dataset sequences in ORB-SLAM3. Each map undergoes full (global) visual-inertial bundle adjustment. For the baseline, we compare our solver to the existing LDLT implementation based on g2o, using the original LM algorithm without modified termination criteria [1] for more consistent behaviour. We also include a PCG baseline, since our implementation uses PCG. For all solvers, the parameter recovery step is skipped so that all optimizations run on the same map. The optimization is set to run for a maximum of 100 iterations. Block ordering is enabled for LDLT as we found this improves the runtime performance. For g2o-PCG, we use a maximum of 50 iterations, with $1\times 10^{-6}$ absolute tolerance, while for our solver, we use a maximum of 100 iterations, with $1\times 10^{-12}$ absolute tolerance, and a rejection ratio of 10. The reason for different termination criteria is because g2o computes and solves a reduced system, instead of using the full Hessian, and we determined that this requires stricter stopping criteria for Graphite to show similar convergence. Additionally, both solvers use block Jacobi preconditioning.
Results: The results for TUM-VI are given in Table IV. Note that there are approximately four vertices created for each keyframe, due to the inclusion of velocity and bias variables (e.g. a map with 3000 keyframes will have 12000 pose vertices). The LDLT baseline converges to the lowest $\chi^{2}$ error for all maps due to computing the exact parameter step at each iteration, at the cost of extremely high runtimes. Compared to the PCG baseline, our solver achieves a relative speedup of up to $59\times$ on the desktop and up to $4\times$ on the Orin Nano, despite using the full Hessian. This difference in speedup is in part due to the Orin Nano having far fewer CUDA cores and operating at lower clock speeds than its desktop counterpart (Table I). Moreover, while lower precision configurations take less time on the Jetson, double precision configurations generally achieve lower $\chi^{2}$ values and run for more iterations. Also, the Orin Nano reports higher memory usages due to the different memory architecture, so we also report the usages minus the baseline usage. Overall, Graphite achieves a significant reduction in time, enabling faster large-scale visual-inertial optimization.
VI Conclusion
We have implemented a general GPU-accelerated non-linear optimization framework based on Levenberg-Marquardt for sparse estimation problems in CUDA C++, which enables the sharing of data types and functions between real-time applications and optimization tasks. Graphite achieves a relatively lower memory footprint by supporting in-place optimization, inferring matrix structures from the optimizable graph, supporting mixed-precision solving, and enabling matrix-free methods. Our results show that our solver enables large-scale optimization over complex constraints for desktop and resource-constrained devices.
Discussion: Currently, Graphite only supports Levenberg-Marquardt and PCG; direct solvers are not yet supported, so the computed parameter step is always inexact, although iterative methods scale better and use less memory. It also assumes that all constraints need observations and constraint data, but this is not always the case. Still, the overhead of storing these is mitigated by using a small data type (e.g. a char). Also, the performance could be improved further by extending vertex descriptors to support reduced camera systems, and using vectorized intrinsics and tensor cores.

### v2

@@
 Abstract
-We present Graphite, a GPU-accelerated non-linear graph optimization framework.
+We present Graphite, a GPU-accelerated nonlinear least squares graph optimization framework.
 It provides a CUDA C++ interface to enable the sharing of code between a real-time application, such as a SLAM system, and its optimization tasks.
@@
 We evaluate Graphite on well-known bundle adjustment problems and find that it achieves similar performance to MegBA, a solver specialized for bundle adjustment, while maintaining generality and using less memory.
-We also apply Graphite to global visual-inertial bundle adjustment on maps generated from stereo-inertial SLAM datasets, and observe speed ups of up to $59\times$ compared to a CPU baseline.
-Our results indicate that our solver enables faster large-scale optimization on both desktop and resource-constrained devices.
+We also apply Graphite to global visual-inertial bundle adjustment on maps generated from stereo-inertial SLAM datasets, and observe speed-ups of up to $59\times$ compared to a CPU baseline.
+Our results indicate that our framework enables faster large-scale optimization on both desktop and resource-constrained devices.
 I Introduction
@@
 To date, there have been several efforts to speed up existing optimization libraries by identifying and offloading expensive steps to GPUs.
-More recently, libraries such as Ceres Solver [4] have implemented GPU acceleration for the linear solver step of trust region optimization algorithms.
-Other work based on g2o [3] focuses on efficiently reducing the size of the linear system to be solved [7] and accelerating numerical differentiation for computing Jacobian matrices [8].
-Yet, while it is possible to improve the performance of existing libraries by offloading individual steps to the GPU, it introduces additional overhead from allocating GPU memory and transferring data between steps.
-An alternative approach is to perform the entire optimization on the GPU.
-Some methods use Python [9] or a domain-specific language (DSL) [10, 11] to allow users to write simple descriptions of nonlinear optimization problems which are then used to automatically generate efficient solvers using complex code transformations.
-Others leverage machine learning frameworks to support batched optimization [12] and differentiable optimization [13, 14] for training learning-based tasks such as feature extraction and feature matching.
+More recently, Ceres Solver [1] implemented GPU acceleration for the linear solver step of trust region optimization algorithms.
+Other work based on g2[18] focuses on efficiently reducing the size of the linear system to be solved [10] and accelerating numerical differentiation for computing Jacobian matrices [17].
+Yet, while it is possible to improve the performance of existing libraries by offloading individual steps to the GPU, doing so introduces additional overhead from allocating GPU memory and transferring data between steps.
+In contrast, new GPU-based frameworks have emerged which accelerate various steps of the optimization.
+Some use Python [32] or a domain-specific language (DSL) [7, 19] to allow users to write simple descriptions of nonlinear optimization problems that are used to automatically generate efficient solvers using complex code transformations.
+Others leverage machine learning frameworks to support batched optimization [15] and differentiable optimization [34, 23, 31] for training models for tasks such as feature extraction and feature matching.
 In addition, many solvers are specialized for specific classes of problems such as bundle adjustment [26], achieving large speedups.
@@
 Meanwhile, SLAM systems may represent variables as complex classes which consist of other classes.
-For example, an optimizable pose class may consist of multiple SE(3) transformations and cameras [1].
-Existing solvers cannot represent these classes with a vector of numbers or predetermined types because they cannot model their complex behaviours and data dependencies.
+For example, an optimizable pose class may consist of multiple SE(3) transformations and camera references [3].
+Existing solvers cannot represent these classes with a vector of numbers or predetermined data types because they cannot model their complex behaviours and data dependencies.
 Additionally, researchers develop real-time SLAM systems in a specific language of their choice, e.g., C++, making it difficult to use DSLs or Python-based optimizers, since they require optimizable data types and functionality to be faithfully reimplemented for another language.
@@
 Even worse, embedded platforms often have no dedicated GPU memory, so the CPU and GPU must compete for the same memory.
-In this paper, we present Graphite, a GPU-accelerated nonlinear optimization framework for estimation problems, which implements several techniques for balancing runtime performance and GPU memory usage, to enable large-scale optimization for desktop and embedded scenarios.
+In this paper, we present Graphite, a GPU-accelerated nonlinear least squares graph optimization framework, which implements several techniques for balancing runtime performance and GPU memory usage, to enable large-scale optimization for desktop and embedded scenarios.
 It allows for mixed-precision solving using 64-bit, 32-bit, and 16-bit floating-point precisions, enabling faster and more memory-efficient optimization of graphs.
@@
 In addition, the iterative linear solver is aware of the structure of the graph, bypassing explicit sparse matrix formats and their associated memory costs.
-To support real-time SLAM and odometry, the framework provides a CUDA C++ interface, and uses a batching model which supports in-place optimization, meaning that users can use their own data types, without having to transform them into a solver-compatible format.
+To support real-time SLAM and odometry, the framework provides a CUDA C++ interface, and uses a batching model which supports in-place optimization, allowing users to optimize GPU-accessible data without first transforming it into a solver-specific format.
 This avoids memory usage and runtime overhead from unnecessary copying, as well as explicit data transfers on platforms which share CPU and GPU memory.
-To demonstrate the effectiveness and flexibility of our framework, we evaluate our solver on bundle adjustment problems, and find it performs comparably to MegBA [15] while using up to 78% less GPU memory.
-We also reimplement visual-inertial bundle adjustment on the GPU inside ORB-SLAM3, which can take hundreds of seconds on the CPU for sufficiently large problems, and demonstrate a speed up of up to $59\times$ on maps generated by processing stereo-inertial SLAM datasets [17], while using under 1 GiB of GPU memory on the desktop.
+To demonstrate the effectiveness and flexibility of our framework, we evaluate it on bundle adjustment problems, and find it performs comparably to MegBA [26] while using up to 78% less GPU memory.
+We also reimplement visual-inertial bundle adjustment on the GPU inside ORB-SLAM3, which can take hundreds of seconds on the CPU for sufficiently large problems, and demonstrate a speed up of up to $59\times$ on maps generated by processing stereo-inertial SLAM datasets [28], while using under 1 GiB of GPU memory on a desktop machine.
 The code is available at https://github.com/sfu-rsl/graphite.
@@
   • We apply our framework to global visual-inertial bundle adjustment inside a SLAM system, which consists of 7 types of constraints and 5 types of variables.
-  • We evaluate our solver across different precisions and differentiation modes, using well-known bundle adjustment and SLAM datasets.
+  • We evaluate our framework across different precisions and differentiation modes, using well-known bundle adjustment and SLAM datasets.
 II Background
@@
 Each constraint represents a relationship between estimated parameters, an observation, and the error.
-The goal is then to find the parameters which minimize the total error across all constraints ($\mathcal{C}$).
+The goal is then to find the parameters that minimize the total error across all constraints ($\mathcal{C}$).
 This may be formulated as
@@
 These problems may be expressed using an optimization framework [6, 18, 1] and solved using algorithms such as Powell’s dog leg or Levenberg-Marquardt.
-Levenberg-Marquardt for Nonlinear Least Squares: Given the non-linear optimization problem in Eq.
+Levenberg-Marquardt for Nonlinear Least Squares: Given the nonlinear optimization problem in Eq.
 1, the Levenberg-Marquardt algorithm (LM) can be used to iteratively find new parameters to drive the error towards a local minimum.
-LM combines aspects of Gauss-Newton and gradient descent to achieve better convergence with the use of a damping factor.
+LM modifies Gauss-Newton using a damping factor $\lambda$ to achieve better convergence.
 It exploits the first-order Taylor expansion
@@
 […]
-$\footnotesize(J^{\top}\Sigma^{-1}J+\lambda I)\Delta x=-J^{\top}\Sigma^{-1}r.$ (5)
-This can be more simply rewritten in terms of the regularized Hessian approximation $H$ and gradient $b$ as
-$H\Delta x=-b.$ (6)
-Solving this linear system in each iteration yields $\Delta x$, which represents a candidate parameter step to reduce the total error. Here, $H$ is a sparse positive semi-definite block matrix.
-Visual-Inertial Optimization: Given a set of keyframes $\kappa$ and $l$ map points, the goal of visual bundle adjustment is to find new parameters which better satisfy camera reprojection error constraints.
+$\footnotesize(J^{\top}\Sigma^{-1}J+\lambda D^{\top}D)\Delta x=-J^{\top}\Sigma^{-1}r.$ (5)
+This can be more simply rewritten in terms of the regularized Hessian approximation $H_{\lambda}$ and gradient $b$ as
+$\footnotesize H_{\lambda}\Delta x=-b.$ (6)
+Solving this linear system in each iteration yields $\Delta x$, which is a candidate parameter step to reduce the total error.
+Here, $H_{\lambda}$ is a sparse symmetric positive definite block matrix [2].
+Visual-Inertial Optimization: Given a set of keyframes $\kappa$ and $l$ map points, the goal of visual bundle adjustment is to find new parameters that better satisfy camera reprojection error constraints.
 The cost may be formulated as Eq.
 7  [3]
-$\footnotesize\sum_{j=0}^{l-1}\sum_{i\in\kappa^{j}}\rho_{\text{Huber}}(\|r_{ij}\|_{\Sigma_{ij}^{-1}}^{2})$ (7)
-where $r_{ij}$ is the residual vector of the reprojection constraint between keyframe $i$ and observed map point $j$.
-In other words, $r_{ij}$ is the difference, in 2D coordinates, between the measured observation of a map point in a keyframe image and its modelled observation computed from the camera’s projection function using current keyframe and map point parameters.
+$\footnotesize\sum_{j=0}^{l-1}\sum_{i\in\kappa^{j}}\rho_{\text{Huber}}(\|r_{ij}\|_{\Sigma_{ij}}^{2})$ (7)
+where $r_{ij}$ is the residual vector of the reprojection constraint between observed map point $j$ and keyframe $i$.
+In other words, $r_{ij}$ is the difference, in 2D coordinates, between the computed camera projection of a map point in a keyframe image using current parameters and its measured observation.
 To reduce sensitivity to feature mismatches between keyframes and potential outliers, a robust Huber kernel $\rho_{\text{Huber}}$ is used [3].
@@
 For k+1 keyframes, this is represented as [3]
-$\footnotesize\sum_{i=1}^{k}\|r_{I_{i-1,i}}\|_{\Sigma_{I_{i,i+1}}^{-1}}^{2}+\sum_{i=1}^{k}\|r_{b_{i-1,i}}\|_{\Sigma_{b_{i,i+1}}^{-1}}^{2}$ (8)
+$\footnotesize\sum_{i=1}^{k}\|r_{I_{i-1,i}}\|_{\Sigma_{I_{i-1,i}}}^{2}+\sum_{i=1}^{k}\|r_{b_{i-1,i}}\|_{\Sigma_{b_{i-1,i}}}^{2}$ (8)
 where $r_{I_{i-1,i}}$ is the inertial residual between consecutive keyframes, representing the discrepancy between preintegrated IMU measurements and the change computed from keyframe parameters.
 These parameters include rotation, velocity, and position.
-Similarly, $r_{b_{i-1,i}}$ is the accelerometer and gyroscope bias change between consecutive keyframes.
+Similarly, $r_{b_{i-1,i}}$ is the accelerometer and gyroscope bias residual between consecutive keyframes.
 GPU Programming: Modern GPUs are based on massively parallel Single-Instruction, Multiple-Thread (SIMT) architecture, where threads execute common instructions in groups.
-CUDA [18] is a programming model which enables writing specialized functions, called kernels, which can be dispatched onto a GPU in order to processes a batch of similar tasks in parallel.
+CUDA [22] is a programming model that enables writing specialized functions, called kernels, which can be dispatched onto a GPU in order to process a batch of similar tasks in parallel.
 CUDA uses a hierarchical execution model where threads are grouped into 3D thread blocks.
@@
 As our work proposes a GPU-based approach for non-linear least squares optimization for problems such as bundle adjustment, we review literature on these topics.
-Although there are several versatile libraries for least squares optimization in robotics and computer vision [2, 3, 4, 19, 20, 21], they are primarily CPU-based and they generally limit GPU acceleration to the linear solver stage [4].
-Our framework differs by performing the entire optimization on the GPU, to enable greater parallelization for variable and constraint operations, and avoid data transfers between intermediate steps to improve performance.
-There are also a number of GPU-accelerated solvers based on domain-specific languages (DSLs) [9, 10, 11] or Python-based machine learning frameworks [12, 13, 14].
+Although there are several versatile libraries for least squares optimization in robotics and computer vision [6, 18, 1, 27, 29, 20], they are primarily CPU-based and generally limit GPU acceleration to the linear solver stage [1].
+Our framework differs by performing the entire optimization on the GPU, to enable greater parallelization for variable and constraint operations and avoid data transfers between intermediate steps to improve performance.
+There are also a number of GPU-accelerated solvers based on domain-specific languages (DSLs) [32, 7, 19] or Python machine learning frameworks [15, 34, 23, 31].
 However, these approaches may result in code duplication, lack the expressiveness needed to model complex types, or involve explicit language interoperation, depending on the application.
-Our framework avoids this by allowing problems to be described directly using user-defined types, preventing code duplication, and enabling easier prototyping.
+Our framework avoids this by allowing problems to be described directly using user-defined data types, preventing code duplication, and enabling easier prototyping.
 Several prior works also explore specialized implementations for bundle adjustment on the GPU, applying direct or iterative solving techniques [26, 33, 14, 9, 4, 8, 13].
-In practice, these libraries are designed and implemented for a specific type of visual bundle adjustment which only uses camera poses, 3D points, and binary constraints.
-Meanwhile, SLAM and odometry systems often add several new constraints by incorporating inertial information into the optimization [1, 28], where some constraints have as many as six variables each.
-At the same time, they may perform other optimization tasks such as motion-only bundle adjustment and pose graph optimization [1, 28].
-Our framework resolves this limitation by supporting user-defined types along with unary, binary, and n-ary GPU constraints.
+While performant, these works target a specific type of visual bundle adjustment which only uses camera poses, 3D points, and binary constraints.
+Meanwhile, SLAM and odometry systems often add several new constraints by incorporating inertial information [3, 25], where some constraints have as many as six variables.
+They may also perform other optimization tasks such as motion-only bundle adjustment and pose graph optimization [3, 25].
+Our framework resolves this limitation by supporting user-defined data types along with unary, binary, and n-ary GPU constraints.
 While prior GPU-accelerated solvers utilize code generation and leverage existing machine learning frameworks, we adopt a design inspired by libraries such as g2o and Ceres Solver, to support the level of functionality needed for visual-inertial SLAM [3] and odometry systems [25], discussed further in Section V.
@@
 IV Design
-Our optimization framework, based on the Levenberg-Marquardt algorithm (Section II), allows for an optimizable graph to be constructed from user-defined types and unary, binary, and n-ary constraints, to support a wide variety of optimization problems.
-Its design is guided by the need to minimize the amount of time needed for optimization while also using as little GPU memory as possible, which is crucial for resource-constrained systems.
-Next, we describe aspects of our framework that are relevant to this approach.
-IV-A Overview
-Graphite implements graph-based optimization, where all major steps run on the GPU.
-It allows optimizable variables (vertices) to be represented by classes defined by the user, which avoids the need to first transform variables into a specific format for optimization, a step which is often required by other libraries.
-These variables can then be iteratively refined in-place on the GPU, using a batching method described in Section IV-B.
-As is standard with other frameworks, Graphite supports fast analytic differentiation (Section IV-C) and convenient automatic differentiation (Section IV-D) for computing Jacobian matrices used by the optimization algorithm.
-To perform the optimization process quickly and in a memory-efficient manner, Graphite also supports mixed-precision solving (Section IV-E).
-IV-B Batching
-Suppose a user wants to optimize noisy 2D points so that they lie on the circumference of a circle.
+Figure 1: An example Graphite workflow.
+The user first defines properties for their variables and constraints, then constructs a graph by creating descriptors, and finally optimizes their variables using Levenberg-Marquardt.
+Graphite implements nonlinear least squares graph optimization on the GPU, based on the Levenberg-Marquardt algorithm (Section II).
+Its design is guided by the need to minimize the amount of time required for optimization while also using as little GPU memory as possible, which is crucial for resource-constrained systems.
+An optimization problem is represented by a graph constructed from a set of descriptors, which are collections of optimizable vertices (variables) and unary, binary, and n-ary constraints.
+Descriptors provide a batching mechanism for the GPU and allow vertices and constraints to be represented with user data types that are compatible with CUDA device code (Section IV-A).
+This avoids the need to first transform variables into a solver-specific format for optimization, a step which is often required by other libraries.
+As is standard with other frameworks, Graphite supports fast analytic differentiation (Section IV-B) and convenient automatic differentiation (Section IV-C) for computing Jacobian matrices used by the optimization algorithm.
+To perform the optimization process quickly and in a memory-efficient manner, Graphite also supports mixed-precision solving (Section IV-D).
+We show a running example of the workflow in Figure 1 for a 2D point optimization problem, where a user adjusts 2D points stored in GPU-accessible memory to lie on the boundary of a circle.
+They first define static properties for their 2D points and circle equation constraints.
+Next, they create the corresponding descriptors and construct a graph.
+Lastly, the user passes the graph to the Levenberg-Marquardt algorithm along with desired options, optimizing the points in-place.
+We discuss this process with concrete examples next.
+IV-A Descriptor Batching Model
 A typical approach to offloading work onto a GPU is to organize similar tasks into batches, and process tasks within a batch concurrently across numerous threads in order to take advantage of the massively parallel architecture.
 However, existing graph-based optimizers store different types of vertices and constraints indiscriminately together in the graph.
-This is not ideal for batching work on the GPU, because each thread processing an item (a vertex or constraint) may execute a different code path, which is detrimental to performance [29].
+This is not ideal for batching work on the GPU, because each thread processing an item (a vertex or constraint) may branch and execute a different code path, which is detrimental to performance [22].
 Therefore, to enable a GPU-friendly approach for graph-based optimization, Graphite allows users to construct an optimizable graph from collections of vertices and constraints of the same type.
-This enables GPU threads to process batches of identical items, which is better for performance [29].
-We represent these collections using descriptors, which we describe next.
-We illustrate how the user models the point optimization problem in Graphite, using mixed-precision where the user selects single precision for the descriptors, and 16-bit brain floating point for the linear system.
-First, the user creates a vertex descriptor for a collection of noisy points for the optimization problem, as shown in Figure 1.
-They specify single precision for the descriptors (Line 1) and 16-bit brain floating point for the linear system (Line 2).
-In Lines 4-6, they create a descriptor, reserve memory for the desired number of vertices, and add the descriptor to the graph.
+This enables GPU threads to process batches of identical items, which minimizes branching.
+We represent these collections using descriptors, which we describe next with our point optimization example.
+We demonstrate how the user models the point optimization problem from Figure 1 in Graphite, where the user selects single precision for the variables and constraints, and 16-bit brain floating point for the linear system.
+First, the user defines static properties for their vertices (points) and constraints (based on the circle equation), which are necessary for creating descriptors, as shown in Figure 1.
+At compile time, these properties are used to specialize each CUDA kernel, so that each thread carries out the same type of computation, improving performance.
+For a vertex descriptor, these properties include information such as the underlying data type (class), the parameter block size, a method to get its parameter block, and a method to update each vertex for a new step.
+1 shows how the user defines the properties for a 2D point descriptor.
+They specify the underlying vertex data type as a $2\times 1$ Eigen [11] matrix (Line 2, Line 7), the length of the parameter block as 2 (Line 6), the method to get the parameter block (Lines 9 to 13), and the update method (Lines 15-18).
+1// Point definition
+2template <typename T> using Point = Eigen::Matrix<T, 2, 1>;
+3
+4// Traits for Point
+5template <typename T, typename S> struct PointTraits {
+6 static constexpr size_t dimension = 2;
+7 using Vertex = Point<T>;
+8
+9 template <typename P>
+10 d_fn static void parameters(const Vertex &vertex, P *parameters) {
+11 Eigen::Map<Eigen::Matrix<P, dimension, 1>> params_map(parameters);
+12 params_map = vertex.template cast<P>();
+13 }
+14
+15 d_fn static void update(Vertex &vertex, const T *delta) {
+16 Eigen::Map<const Eigen::Matrix<T, dimension, 1>> d(delta);
+17 vertex += d;
+18 }
+19};
+20
+21// Type alias
+22template <typename T, typename S>
+23using PointDescriptor = VertexDescriptor<T, S, PointTraits<T, S>>;
+Listing 1: Properties of a vertex descriptor for a 2D point.
+Next, the user defines static properties for their constraints, including which vertex descriptors are involved, the data types (classes) of the observations and any non-optimizable data, the loss function used, and the differentiation mode (discussed in Section IV-B and Section IV-C).
+These properties also specify how to compute the residual for a constraint, and how to compute the Jacobians when using analytic differentiation (Section IV-B).
+2 shows this for the point optimization problem.
+The user specifies a residual of size 1 (Line 3), specifies the vertex descriptor and data types (Lines 4-6), chooses the default loss function (Line 7), and chooses to use automatic differentiation (Line 8), so no Jacobian evaluation function is specified.
+Additionally, Lines 10-16 tell the descriptor how to compute the residual based on the equation of a circle.
+1// Traits for the circle constraint
+2template <typename T, typename S> struct CircleTraits {
+3 static constexpr size_t dimension = 1;
+4 using VertexDescriptors = std::tuple<PointDescriptor<T, S>>;
+5 using Observation = T;
+6 using Data = Empty; // unused - not passed to error function
+7 using Loss = DefaultLoss<T, dimension>;
+8 using Differentiation = DifferentiationMode::Auto;
+9
+10 template <typename D>
+11 d_fn static void error(const D *point, const T &obs, D *error) {
+12 const auto x = point[0]; // Note: A const Point<T>& can also be passed in.
+13 const auto y = point[1]; // The framework automatically determines
+14 const auto r = obs; // how to call the user’s function.
+15 error[0] = x * x + y * y - r * r;
+16 }
+17};
+18
+19template <typename T, typename S>
+20using CircleDescriptor = FactorDescriptor<T, S, CircleTraits<T, S>>;
+Listing 2: Properties for the circle equation constraint descriptor.
+We specify properties of descriptors in this way because it lets users declare the behaviour of their types and automatically generate optimized GPU code, without relying on runtime lookups (as observed in some CPU-based optimizers [18]) to decide how each item in the graph is processed.
+Following Figure 1, after defining these properties, the user creates descriptors in order to construct an optimizable graph.
+The user first creates a vertex descriptor for the collection of points, as shown in 3.
+Before doing so, they declare a graph with the desired precisions in Lines 1-3.
+Then in Lines 4-6, they create the vertex descriptor, reserve memory for the desired number of vertices, and add the descriptor to the graph.
 Then from Lines 8-10, they add a pointer for each vertex to the descriptor, and associate it with an identifier.
-In creating this descriptor, the user defines a batch of vertices which is processed efficiently on the GPU across many threads.
-using FP = float; // vertex, constraint precision
-using SP = __nv_bfloat16; // linear system precision
-[…]
-point_desc.reserve(num_vertices); // Reserve memory for num_vertices
-graph.add_vertex_descriptor(&point_desc); // Add to graph
-[…]
-point_desc.add_vertex(vertex_id, &points[vertex_id]); // Add points
-Figure 1: Creating a vertex descriptor for 2D points.
-Next, they create a descriptor for the constraints in Figure 2.
-As in the previous example, they create a descriptor, reserve memory for the desired number of constraints (one constraint for each vertex), and add the descriptor to the graph (Lines 1-3).
+In creating this descriptor, the user defines a batch of vertices which is processed efficiently across many GPU threads.
+1using FP = float; // vertex, constraint precision
+2using SP = __nv_bfloat16; // linear system precision
+3Graph<FP, SP> graph; // Declare graph
+4auto point_desc = PointDescriptor<FP, SP>(); // Create descriptor
+5point_desc.reserve(num_vertices); // Reserve memory for num_vertices
+6graph.add_descriptor(&point_desc); // Add to graph
+7
+8for (size_t vertex_id = 0; vertex_id < num_vertices; ++vertex_id) {
+9 point_desc.add_vertex(vertex_id, &points[vertex_id]); // Add points
+10}
+Listing 3: Creating a vertex descriptor for 2D points.
+Next, they create a descriptor for the constraints in 4.
+As in the previous example, they initialize the descriptor, reserve memory for the desired number of constraints (one constraint for each vertex), and add the descriptor to the graph (Lines 1-3).
 They initialize the default loss function on Line 4, and then create the desired number of constraints (Lines 6-8).
-When creating each constraint, the user specifies the identifiers of the vertices involved, the observation (in this case, the radius of the circle), the precision matrix (nullptr defaults to the identity matrix), the data for non-optimizable variables (not used), and the loss function parameters.
+When creating each constraint, the user specifies the identifiers of the vertices involved, the observation (the radius of the circle), the precision matrix (nullptr defaults to the identity matrix), optional constraint data (unused), and the loss function parameters.
 Again, by constructing this descriptor, the user enables each constraint to be processed concurrently and efficiently on the GPU.
-auto factor_desc = CircleFactor<FP, SP>(&point_desc); // Create descriptor
-factor_desc.reserve(num_vertices); // Reserve memory for constraints
-graph.add_factor_descriptor(&factor_desc); // Add descriptor to graph
-const auto lf = DefaultLoss<FP, 1>(); // loss function
-[…]
-factor_desc.add_factor({vertex_id}, radius, nullptr, 0, lf);
-Figure 2: Creating a descriptor for constraints.
-At compile time, the properties of each descriptor are used to specialize each CUDA kernel for operations over a specific constraint or vertex type, so that each thread carries out the same type of computation, which improves performance.
-For a vertex descriptor, these properties include information such as the underlying data type (class), the size of its parameter block, a method to get its parameter block, and a method to update each vertex for a new step.
-Figure 3 shows how the user defines the properties for the 2D point descriptor.
-They specify the underyling vertex data type as a $2\times 1$ Eigen [30] matrix (Line 2, Line 7), the length of the parameter block as 2 (Line 6), the method to get the parameter block (Lines 9 to 11), and the update method (Lines 13-16).
-// Point definition
-template <typename T> using Point = Eigen::Matrix<T, 2, 1>;
-[…]
-template <typename T, typename S> struct PointTraits {
-static constexpr size_t dimension = 2;
-using Vertex = Point<T>;
-[…]
-return vector_to_array<T, dimension>(vertex);
-[…]
-Eigen::Map<const Eigen::Matrix<T, dimension, 1>> d(delta);
-vertex += d;
-[…]
-template <typename T, typename S>
-using PointDescriptor = VertexDescriptor<T, S, PointTraits<T, S>>;
-Figure 3: Properties of a vertex descriptor for a 2D point.
-In general, properties for a constraint descriptor include which vertex descriptors are involved, the data types (classes) of the observations and any non-optimizable data, the loss function used, and the differentiation mode (discussed in Section IV-C and Section IV-D).
-These properties also specify how to compute the residual for a constraint, and how to compute the Jacobians when using analytic differentiation (Section IV-C).
-Figure 4 shows this for the point optimization problem.
-The user specifies a residual of size 1 (Line 3), specifies the point vertex descriptor (Line 4), chooses the default loss function (Line 7), and chooses to use automatic differentiation (Line 8), so no Jacobian evaluation function is specified.
-Additionally, Lines 10-18 tell the descriptor how to compute the residual based on the squared distance between the point and the radius.
-We specify properties of descriptors in this way because it lets users declare the behaviour of their types and automatically generate optimized GPU code, without relying on runtime lookups (as observed in some CPU-based optimizers [3]) to decide how each item in the graph is processed.
-// Factor traits for the circle constraint
-template <typename T, typename S> struct CircleFactorTraits {
-static constexpr size_t dimension = 1;
-using VertexDescriptors = std::tuple<PointDescriptor<T, S>>;
-using Observation = T;
-using Data = unsigned char;
-using Loss = DefaultLoss<T, dimension>;
-using Differentiation = DifferentiationMode::Auto;
-[…]
-hd_fn static void error(const D *point, const T *obs, D *error,
-const std::tuple<Point<T> *> &vertices,
-const Data *data) {
-auto x = point[0];
-auto y = point[1];
-auto r = obs[0];
-error[0] = x * x + y * y - r * r;
-[…]
-using CircleFactor = FactorDescriptor<T, S, CircleFactorTraits<T, S>>;
-Figure 4: Properties of a constraint checking a point’s distance from the circumference of a circle with radius $r$.
-After the user finishes adding descriptors to the graph and populating them with information about each vertex and constraint, they pass the graph to an optimization algorithm, along with a linear solver (discussed further in Section IV-E).
-The user optimizes the graph for the point optimization example in Figure 5.
-First, they decide to fix the value of the last point they created (Line 1) and decide to disable the constraint which was created for the third point by changing its level to 1, as in g2o [3] (Line 2).
-Next, the user configures the iterative linear solver, by selecting an identity matrix for the preconditioner (Line 3) and then instantiating the solver (Line 4) with their desired parameters.
-They choose 10 iterations for Levenberg-Marquardt (Line 6), and create one stream (Line 7), since each constraint only has one vertex.
-On Line 8, they set the optimization level to 0, meaning that any constraints with a level higher than 0, such as the constraint on Line 2, will be excluded from the optimization.
-Finally, they perform the optimization by calling Levenberg-Marquardt in Lines 9-10.
-point_desc.set_fixed(num_points - 1, true); // Set the last vertex as fixed
-factor_desc.set_active(2, 0x1); // Disable third constraint for point 2
-IdentityPreconditioner<FP, SP> preconditioner;
-PCGSolver<FP, SP> solver(50, 1e-6, 10.0, &preconditioner);
-[…]
-StreamPool streams(1); // chose 1 stream for 1 vertex per constraint
-constexpr uint8_t optimization_level = 0;
-optimizer::levenberg_marquardt<FP, SP>(&graph, &solver, iterations, 1e-6,
-optimization_level, streams);
-Figure 5: Example of optimizing noisy 2D points.
-IV-C Analytic Differentiation
+1auto circle_desc = CircleDescriptor<FP, SP>(&point_desc); // Create descriptor
+2circle_desc.reserve(num_vertices); // Reserve memory for constraints
+3graph.add_descriptor(&circle_desc); // Add descriptor to graph
+4const auto loss = DefaultLoss<FP, 1>(); // loss function
+5
+6for (size_t vertex_id = 0; vertex_id < num_vertices; ++vertex_id) {
+7 circle_desc.add_factor({vertex_id}, radius, nullptr, Empty(), loss);
+8}
+Listing 4: Creating a descriptor for constraints.
+After the user finishes constructing the graph, they pass it to an optimization algorithm, along with various options, including the desired linear solver (discussed in Section IV-D), optimizing the points in-place (Figure 1).
+For the point optimization example, this is shown in 5.
+First, the user decides to fix the value of the last point they created (Line 1) and decides to disable the constraint which was created for the third point by changing its level to 1, as in g2[18] (Line 2).
+Next, the user configures an iterative linear solver, by selecting an identity matrix for the preconditioner (Line 3) and then instantiating the solver (Line 4) with their desired parameters.
+They create one stream (Line 5), since each constraint only has one vertex.
+Lastly, they configure the Levenberg-Marquardt algorithm on Lines 7-12, and run the optimization on Line 13.
+1point_desc.set_fixed(num_vertices - 1, true); // Set the last vertex as fixed
+2circle_desc.set_active(2, 0x1); // Disable third constraint for point 2
+3IdentityPreconditioner<FP, SP> preconditioner;
+4PCGSolver<FP, SP> solver(50, 1e-6, 10.0, &preconditioner);
+5StreamPool streams(1); // chose 1 stream for 1 vertex per constraint
+6
+7optimizer::LevenbergMarquardtOptions<FP, SP> options;
+8options.solver = &solver;
+9options.initial_damping = 1e-6;
+10options.iterations = 10;
+11options.optimization_level = 0;
+12options.streams = &streams;
+13optimizer::levenberg_marquardt<FP, SP>(&graph, &options);
+Listing 5: Example of optimizing noisy 2D points.
+IV-B Analytic Differentiation
 In graph-based optimization, constraints may involve one or more vertices, each of which generate a Jacobian matrix.
@@
 A typical approach on the CPU is to compute each matrix for a constraint sequentially [18, 1].
-However, to take advantage of the many GPU threads available, Graphite computes each Jacobian concurrently in separate streams (we describe streams in Section II).
+However, to take advantage of the many GPU threads available, Graphite attempts to compute each Jacobian concurrently in separate streams (described in Section II).
 Additionally, Graphite allows these Jacobians to be evaluated dynamically [33], at the point of computation (e.g.
-matrix-vector multiplication), to enable matrix-free solving methods which use less GPU memory.
-IV-D Automatic Differentiation
+matrix-vector multiplication), to enable matrix-free solving methods which use less memory.
+IV-C Automatic Differentiation
 Another method to compute Jacobians is through forward mode automatic differentiation.
-For example, Ceres Solver computes the gradient of a residual function using dual numbers, which consist of a real component $a$, for the computed value, and an infinitesimal component $b\varepsilon$, where $\varepsilon^{2}=0$, for the corresponding partial derivative [4].
+For example, Ceres Solver computes the gradient of a residual function using dual numbers in the form $a+b\varepsilon$, where a value is represented by real component $a$, and the partial derivative is represented by infinitesimal component $b$, where $\varepsilon^{2}=0$ [1].
 To extend this for multiple inputs, Ceres implements a data structure called a Jet, which replaces the infinitesimal component with a vector [1].
@@
 Figure 2: A CUDA thread calculates $\frac{\partial f}{\partial x_{3}}(\vec{x})$ by replacing the input with a vector of dual numbers and writes it to the corresponding Jacobian column.
-IV-E Mixed-Precision Solver
-To enable mixed-precision solving, which enables the optimization to use less memory, Graphite supports the customization of floating point precisions for descriptors and the linear system (Eq.
+IV-D Mixed-Precision Solver
+To enable mixed-precision solving, which allows the optimization to use less memory, Graphite supports the customization of floating point precisions for vertices and the linear system (Eq.
 6).
 Essentially, these two precisions can be customized according to trade-offs between convergence, runtime performance, and memory usage.
-A typical approach would be to use a higher precision (double or single precision) for the descriptors to maintain numerical stability, and an equal or lower precision (single precision or brain floating point), to reduce the amount of memory needed to store the Jacobian matrices [32].
+A typical approach would be to use a higher precision (double or single precision) for the vertices to maintain numerical stability, and an equal or lower precision (single precision or brain floating point) for the linear system to reduce the amount of memory needed to store it [24].
 In contrast, existing solvers support only one precision at a time [26], or only allow casting the linear system to single precision [1].
-As shown in Section V, using lower precisions greatly reduces the overall optimization time and amount of GPU memory required.
+As shown in Section V, using lower precisions greatly reduces the overall optimization time and the amount of GPU memory required.
 As described in Section II, each iteration of Levenberg-Marquardt optimization computes a parameter step $\Delta x$.
@@
 To mitigate this, we adopt techniques which are known to help with solving larger systems.
-These include, clamping values along the Hessian diagonal [4], rescaling the column of each Jacobian block by the square-root inverse of the corresponding Hessian diagonal value [4, 32], and normalizing the PCG residual in each iteration [33].
+These include clamping values along the Hessian diagonal [1], rescaling the columns of each Jacobian block according to the Hessian diagonal [1, 24], and normalizing the PCG residual in each iteration [12].
 In applying PCG, to improve convergence, we also use a block Jacobi preconditioner [2], because it is relatively simple to compute and does not require much memory due to its sparse block-diagonal matrix structure.
 We have observed that storing this preconditioner at too low of a precision can also result in convergence issues.
-Thus, when a 16-bit precision type such as bfloat16 is specified for the linear system, we compute and store this preconditioner at single or double precision, according to the precision of the graph.
+Thus, when a 16-bit precision type such as bfloat16 is specified for the linear system, we compute and store this preconditioner at single or double precision, according to the precision of the variables.
 Additionally, since the PCG solver does not require storing the Hessian matrix, we avoid explicitly constructing sparse matrices for the optimization problem.
 Instead, when performing an operation such as matrix-vector multiplication, the kernel uses the structure of the graph itself.
-This eliminates additional memory and runtime costs associated with converting and storing matrices in different formats, although each kernel must now be aware of the graph structure and whether a variable is fixed or constraint is inactive.
+This eliminates additional memory and runtime costs associated with converting and storing matrices in different formats, although each kernel must now be aware of the graph structure and whether a variable is fixed or a constraint is inactive.
 V Evaluation
-Machine Specs
-12-core Intel Core i7-12700K CPU @ 4.9 GHz
-Desktop 10496-core NVIDIA RTX 3090 GPU @ 1.9 GHz
-64 GB RAM
-6-core ARM Cortex-A78AE CPU @ 1.7 GHz
-Orin Nano 1024-core NVIDIA Ampere GPU @ 1.1 GHz
-8 GB RAM
-TABLE I: Evaluation machine specifications.
-Problem Metric FP64 FP32 FP32-BF16 Ceres-CGNR MegBA-FP64 MegBA-FP32
-MSE 0.43 0.43 0.43 0.43 0.43 0.43
-Dubrovnik-16 Time 0.22 0.11 0.12 1.64 0.17 0.16
-GPU Memory 298 282 278 336 406 352
-MSE 1.67 1.67 1.68 1.67 1.67 1.67
-Trafalgar-21 Time 0.11 0.069 0.068 0.82 0.098 0.088
-GPU Memory 280 274 270 318 326 298
-MSE 0.85 0.85 0.85 0.84 0.84 0.84
-Ladybug-49 Time 0.063 0.041 0.046 0.66 0.086 0.074
-GPU Memory 278 274 270 316 326 300
-MSE 0.67 0.67 0.67 0.67 0.66 0.67
-Venice-1778 Time 6.00 2.80 2.87 43.71 6.96 6.15
-GPU Memory 1944 1190 950 2160 7810 4528
-MSE 1.13 1.13 1.14 1.14 1.12 1.13
-Final-4585 Time 9.62 4.69 3.75 73.49 13.19 9.82
-GPU Memory 3102 1822 1386 3546 13882 7960
-MSE 1.50 1.50 1.50 1.51 OOM OOM
-Final-13682 Time 26.60 11.65 11.67 210.64 OOM OOM
-GPU Memory 9298 5212 3828 10640 OOM OOM
-TABLE II: Time (seconds), mean squared error, and GPU memory usage (MiB) for BAL experiments using different precisions (desktop).
+[b]
+TABLE I: Bundle adjustment results for BAL (desktop), including mean squared error, time (seconds), and GPU memory (MiB).
+BAL Problem Evaluation Graphite Graphite Graphite Ceres MegBA MegBA DeepLM
+Problem Size1 Metric FP64 FP32 FP32-BF16 CGNR FP64 FP32 PCG
+16 Img MSE 0.43 0.43 0.43 0.43 0.43 0.43 0.43
+Dubrovnik-16 22106 Pts Time 0.22 0.11 0.12 1.64 0.17 0.16 3.88
+83718 Obs GPU Memory 298 282 278 336 406 352 814
+21 Img MSE 1.67 1.67 1.68 1.67 1.67 1.67 1.67
+Trafalgar-21 11315 Pts Time 0.11 0.069 0.068 0.82 0.098 0.088 1.61
+36455 Obs GPU Memory 280 274 270 318 326 298 764
+49 Img MSE 0.85 0.85 0.85 0.84 0.84 0.84 0.84
+Ladybug-49 7776 Pts Time 0.063 0.041 0.046 0.66 0.086 0.074 1.14
+31843 Obs GPU Memory 278 274 270 316 326 300 740
+1778 Img MSE 0.67 0.67 0.67 0.67 0.66 0.67 0.67
+Venice-1778 993923 Pts Time 6.00 2.80 2.87 43.71 6.96 6.15 41.67
+5001946 Obs GPU Memory 1944 1190 950 2160 7810 4528 5708
+4585 Img MSE 1.13 1.13 1.14 1.14 1.12 1.13 1.13
+Final-4585 1324582 Pts Time 9.62 4.69 3.75 73.49 13.19 9.82 43.11
+9125125 Obs GPU Memory 3102 1822 1386 3546 13882 7960 7158
+13682 Img MSE 1.50 1.50 1.50 1.51 OOM OOM 1.51
+Final-13682 4456117 Pts Time 26.60 11.65 11.67 210.64 OOM OOM 55.62
+28987644 Obs GPU Memory 9298 5212 3828 10640 OOM OOM 15044
+  • 1
+    Shows the number of camera images (Img), points (Pts), and observations (Obs).
+TABLE II: Bundle adjustment across differentiation modes, including mean squared error, time (seconds), and GPU memory (MiB).
 Problem Metric Analytic Dynamic Auto
 MSE 1.50 1.50 1.50
-Final-13682 Time 26.57 417 38.81
-(FP64) GPU Memory 9298 3988 9298
+Final-13682 Time 26.60 417 38.81
+FP64 GPU Memory 9298 3988 9298
 MSE 1.50 1.50 1.50
-Final-13682 Time 10.97 25.22 11.46
-(FP32) GPU Memory 5212 2556 5212
+Final-13682 Time 11.65 25.22 11.46
+FP32 GPU Memory 5212 2556 5212
 MSE 1.50 1.50 1.50
-Final-13682 Time 10.64 24.48 11.27
-(FP32-BF16) GPU Memory 3828 2500 3828
-TABLE III: Time (seconds), mean squared error, and GPU memory usage (MiB) for across differentiation modes (desktop).
-Sequence Initial $\chi^{2}$ Graph Size Metric g2o-LDLT g2o-PCG FP64 FP32 FP32-BF16
+Final-13682 Time 11.67 24.48 11.27
+FP32-BF16 GPU Memory 3828 2500 3828
+[b]
+TABLE III: Full-inertial bundle adjustment on ORB-SLAM3 maps, including the total $\chi^{2}$ error, time (seconds) and GPU memory (MiB).
+TUM-VI Initial $\chi^{2}$ Graph Evaluation g2o g2o Graphite Graphite Graphite
+Sequence Error Size1 Metric2 LDLT PCG FP64 FP32 FP32-BF16
 3131 KF Final $\chi^{2}$ $3.0918\times 10^{5}$ $4.4965\times 10^{5}$ $4.3850\times 10^{5}$ $4.6842\times 10^{5}$ $5.9817\times 10^{5}$
 outdoors1 $1.0803\times 10^{6}$ 58886 MP Time 1424.40 224.08 4.96 1.71 1.18
-(Desktop) 348548 Co GPU Mem. 678 580 542
+Desktop 348548 Co GPU Mem. 678 580 542
 2032 KF Final $\chi^{2}$ $2.9105\times 10^{5}$ $3.5100\times 10^{5}$ $3.5012\times 10^{5}$ $3.4090\times 10^{5}$ $4.8966\times 10^{5}$
 outdoors2 $7.3020\times 10^{5}$ 57641 MP Time 698.68 147.57 2.80 3.56 1.10
-(Desktop) 338364 Co GPU Mem. 656 564 530
+Desktop 338364 Co GPU Mem. 656 564 530
 1595 KF Final $\chi^{2}$ $2.5617\times 10^{5}$ $3.0916\times 10^{5}$ $3.0354\times 10^{5}$ $0684\times 10^{5} 3.3$$3927\times 10^{5}$
 outdoors3 $6.4905\times 10^{5}$ 36536 MP Time 2957.40 99.69 1.69 1.32 1.13
-(Desktop) 272307 Co GPU Mem. 636 562 534
+Desktop 272307 Co GPU Mem. 636 562 534
 2483 KF Final $\chi^{2}$ $2.2494\times 10^{5}$ $2.9363\times 10^{5}$ $2.9112\times 10^{5}$ $2.9348\times 10^{5}$ $3.6397\times 10^{5}$
 outdoors1 $7.1589\times 10^{5}$ 48621 MP Time 2411.40 512.58 122.61 93.10 27.90
-(Orin Nano) 278661 Co GPU Mem. 5673, 306 5462, 95 5433, 66
+Orin Nano 278661 Co GPU Mem. 5673, 306 5462, 95 5433, 66
 1972 KF Final $\chi^{2}$ $2.1432\times 10^{5}$ $3585\times 10^{5} 3.3$$3731\times 10^{5}$ $3.5705\times 10^{5}$ $3.5889\times 10^{5}$
 outdoors2 $7.3551\times 10^{5}$ 33831 MP Time 1141.40 345.46 53.63 32.58 29.73
-(Orin Nano) 213125 Co GPU Mem. 5128, 328 5031, 231 5017, 217
+Orin Nano 213125 Co GPU Mem. 5128, 328 5031, 231 5017, 217
 1589 KF Final $\chi^{2}$ $1.6574\times 10^{5}$ $1.9508\times 10^{5}$ $1.9439\times 10^{5}$ $2.0736\times 10^{5}$ $2.1015\times 10^{5}$
 outdoors3 $5.3520\times 10^{5}$ 24524 MP Time 2166.1 355.34 93.71 33.53 29.64
-(Orin Nano) 174070 Co GPU Mem. 4674, 282 4473, 81 4449, 57
-TABLE IV: Results of full-inertial bundle adjustment on generated maps. Time given in seconds and GPU memory in MiB.
+Orin Nano 174070 Co GPU Mem. 4674, 282 4473, 81 4449, 57
+  • 1
+    Shows the number of keyframes (KF), map points (MP), and constraints (Co). Each keyframe generates four variables (pose, velocity, IMU biases).
+  • 2
+    For the Orin Nano, we report the peak GPU memory usage and peak usage minus the baseline usage before Graphite runs.
 We evaluate Graphite in terms of the error reduction, runtime performance, and GPU memory usage.
-First, we compare the solver in different precision modes, using problems from Bundle Adjustment in the Large (BAL) [34], a popular dataset for least squares and bundle adjustment solvers.
+First, we compare the solver in different precision modes, using small to large problems from Bundle Adjustment in the Large (BAL) [2], which serves as a common benchmark across optimization libraries.
 Next, we compare analytic, dynamic, and automatic differentiation modes.
-Lastly, we run full-inertial bundle adjustment with Graphite inside ORB-SLAM3 using maps generated from outdoors sequences in TUM-VI, a well-known dataset for evaluating visual-inertial SLAM systems.
-We run these experiments using a desktop machine equipped with a discrete graphics card, and an NVIDIA Jetson Orin Nano Developer Kit, an embedded robotics platform, to assess the performance of Graphite in high-performance and resource-constrained environments (specifications in Table I).
+Lastly, we run full-inertial bundle adjustment with Graphite inside ORB-SLAM3, using maps generated from outdoors sequences in TUM-VI [28], a well-known dataset for evaluating visual-inertial SLAM systems.
+To assess Graphite in high-performance and resource-constrained environments, we run these experiments on a desktop machine with a 12-core Intel Core i7-12700K CPU at 3.8-4.9 GHz, 10496-core NVIDIA RTX 3090 GPU at 1.9 GHz, and 64 GB RAM, and also on an NVIDIA Jetson Orin Nano with a 6-core ARM Cortex-A78AE CPU at 1.7 GHz, 1024-core NVIDIA Ampere GPU at 1.0 GHz, and 8 GB RAM.
+We use unified memory to store optimizable variables.
 V-A Mixed-Precision Modes
@@
 We use bfloat16 rather than 16-bit half precision, because it can approximately represent the same range as single precision.
-For comparison, we also include two GPU-accelerated solvers, Ceres Solver [4], a versatile solver commonly used for SLAM, configured for CUDA CGNR (conjugate gradients on the normal equations), which like our solver, also uses implicit Hessian evaluation, and MegBA [15], a high-performance solver for bundle adjustment with support for single and double precision solving.
+For comparison, we select two general GPU-accelerated optimization libraries with BAL support and similar conjugate gradient solvers, Ceres Solver [1] and DeepLM [15], as well as MegBA [26], a fully-accelerated, high-performance bundle adjuster with single and double precision modes.
 Where available, we use analytic Jacobians, which have the least runtime overhead.
-We generate analytic Jacobians for our solver using wrenfold [31] for symbolic differentiation.
+We generate Jacobian expressions for Graphite using wrenfold [5].
 To allow each optimizer sufficient opportunity to converge, we use a maximum of 50 iterations.
-For PCG in MegBA, the maximum number of iterations is set to 100, while 10 is used for the other configurations, since they solve the full system rather than a reduced one, which involves more work per iteration.
-As shown in Table II, the solver is able to use substantially less memory when using BF16 to store the Jacobian matrices, at the cost of slightly worse convergence.
+For PCG in MegBA, the maximum number of iterations is set to 100, while 10 is used for the other configurations, since they solve the full Hessian system rather than a reduced one, which involves more work per iteration.
+As shown in Table I for single run experiments, Graphite uses substantially less memory when using BF16 to store the Jacobian matrices, at the cost of slightly worse convergence.
 For runtime, FP64 takes the longest, while FP32 and FP32-BF16 take a comparable amount of time.
-Compared to Ceres-CGNR, Graphite takes less time since it accelerates constraint-specific calculations as well.
+Compared to similar libraries, Graphite takes less time since it accelerates constraint-specific calculations as well.
 Although MegBA achieves slightly lower MSE in a few instances, Graphite uses $\frac{1}{4}$ the memory for large problems, since it avoids storing Hessian matrices, and receives a greater speedup when using single precision.
@@
 V-B Differentiation Modes
-The performance of our solver on the desktop using analytic, dynamic, and automatic differentiation modes is given in Table III for Final-13682, the largest dataset in BAL.
-Automatic differentiation allocates the same amount of memory as analytic differentiation, taking marginally more time at lower precisions, and significantly more at FP64.
+We report Graphite’s performance using analytic, dynamic, and automatic differentiation modes in Table II for Final-13682, the largest dataset in BAL.
+Automatic differentiation allocates the same amount of memory as analytic differentiation, taking approximately the same time at lower precisions, and significantly more at FP64.
 Meanwhile, the dynamic mode requires the least amount of GPU memory since Jacobians are not stored, although precision matrices are still used.
@@
 We apply Graphite to full-inertial bundle adjustment in ORB-SLAM3, which introduces several additional variables and constraints to visual bundle adjustment (Section II).
-Implementation: To use Graphite for visual-inertial bundle adjustment, it is necessary to reimplement the vertices and constraints involved in the optimization.
+Implementation: To use Graphite for visual-inertial bundle adjustment, we reimplement vertices and constraints involved in the optimization, external to the library.
 The original data types in ORB-SLAM3 cannot be used as-is, since some contain extraneous data, involve dynamic memory allocations which must be manually transferred from the CPU to the GPU, or implement methods which are not suitable for GPU execution (e.g.
@@
 The remaining vertex data types for the map points, velocities, gyro biases, and accelerometer biases are $3\times 1$ vectors, so we directly represent these using Eigen [11].
-To use these data types with Graphite, we create five vertex descriptors.
+To use these data types with Graphite, we define descriptors, as in 1.
 To model the optimization constraints, we create eight descriptors for monocular and stereo reprojection errors, the inertial residual (with and without Huber loss), gyro and accelerometer bias residuals, and bias priors.
-The residual and analytic Jacobian expressions for each constraint are mostly a direct translation of the original ORB-SLAM3 implementation, although we found that some functionality does not behave as expected on the GPU.
-For example, ORB-SLAM3 normalizes a rotation matrix $R$ by first computing its SVD decomposition and then computing $R=UV^{\intercal}$, but this fails when done in a kernel, so we replace it with a Newton-Schulz iteration.
-For Lie groups, we use Sophus [35], as in the original implementation.
-When constructing the graph, to avoid overhead from reallocating memory, we predetermine the size of the optimization, reserve memory for each descriptor, and preallocate unified memory to store the optimizable variables.
-Experiment: Next, large visual-inertial maps are generated by processing TUM-VI dataset sequences in ORB-SLAM3.
-Each map undergoes full (global) visual-inertial bundle adjustment.
-For the baseline, we compare our solver to the existing LDLT implementation based on g2o, using the original LM algorithm without modified termination criteria [1] for more consistent behaviour.
+Following the original implementation, we use the same expressions for computing the residuals and Jacobians of each constraint along with Sophus [30] for Lie groups.
+When constructing the graph, to avoid overhead from reallocating memory, we predetermine the size of the optimization and reserve the required memory for descriptors and optimizable variables.
+Experiment: Next, we evaluate the performance of visual-inertial bundle adjustment inside ORB-SLAM3 using Graphite.
+Among compatible datasets, we choose outdoors sequences from the TUM-VI dataset, since they generate large maps consisting of thousands of keyframes.
+Within ORB-SLAM3, we perform full (global) visual-inertial bundle adjustment on each of the generated maps.
+For the baseline, we compare Graphite to the existing LDLT-based g2o implementation, using the original LM algorithm without modified termination criteria [3] for more consistent behaviour.
 We also include a PCG baseline, since our implementation uses PCG.
+Note that the challenges discussed in Section III restrict which solvers can be used.
+Limiting factors include language compatibility and sufficient support for user data types involved in complex constraints (e.g.
+with up to six variables).
 For all solvers, the parameter recovery step is skipped so that all optimizations run on the same map.
-The optimization is set to run for a maximum of 100 iterations.
+The optimization is set to run for up to 100 iterations.
 Block ordering is enabled for LDLT as we found this improves the runtime performance.
-For g2o-PCG, we use a maximum of 50 iterations, with $1\times 10^{-6}$ absolute tolerance, while for our solver, we use a maximum of 100 iterations, with $1\times 10^{-12}$ absolute tolerance, and a rejection ratio of 10.
-The reason for different termination criteria is because g2o computes and solves a reduced system, instead of using the full Hessian, and we determined that this requires stricter stopping criteria for Graphite to show similar convergence.
-Additionally, both solvers use block Jacobi preconditioning.
-Results: The results for TUM-VI are given in Table IV.
-Note that there are approximately four vertices created for each keyframe, due to the inclusion of velocity and bias variables (e.g.
-a map with 3000 keyframes will have 12000 pose vertices).
+For g2o PCG, we use a maximum of 50 iterations, with $1\times 10^{-6}$ absolute tolerance, while for our solver, we use a maximum of 100 iterations, with $1\times 10^{-12}$ absolute tolerance, and a rejection ratio of 10.
+The reason for different termination criteria is that g2o computes and solves a reduced system, while Graphite uses the full Hessian and gradient.
+Both solvers apply block Jacobi preconditioning.
+Results: The results for TUM-VI are given in Table III.
 The LDLT baseline converges to the lowest $\chi^{2}$ error for all maps due to computing the exact parameter step at each iteration, at the cost of extremely high runtimes.
-Compared to the PCG baseline, our solver achieves a relative speedup of up to $59\times$ on the desktop and up to $4\times$ on the Orin Nano, despite using the full Hessian.
-This difference in speedup is in part due to the Orin Nano having far fewer CUDA cores and operating at lower clock speeds than its desktop counterpart (Table I).
-Moreover, while lower precision configurations take less time on the Jetson, double precision configurations generally achieve lower $\chi^{2}$ values and run for more iterations.
-Also, the Orin Nano reports higher memory usages due to the different memory architecture, so we also report the usages minus the baseline usage.
+Compared to the PCG baseline, we achieve a relative speedup of up to $59\times$ on the desktop and up to $6\times$ on the Orin Nano, despite using the full Hessian.
+This difference in speedup is in part due to the Orin Nano having far fewer CUDA cores and operating at lower clock speeds than its desktop counterpart.
+Moreover, while lower precision modes take less time on the Jetson, FP64 generally achieves lower $\chi^{2}$ values and runs for more iterations.
+Note that the Jetson reports higher memory usages due to the different architecture, so we also include the peak usage minus the baseline usage before Graphite first runs.
 Overall, Graphite achieves a significant reduction in time, enabling faster large-scale visual-inertial optimization.
 VI Conclusion
-We have implemented a general GPU-accelerated non-linear optimization framework based on Levenberg-Marquardt for sparse estimation problems in CUDA C++, which enables the sharing of data types and functions between real-time applications and optimization tasks.
+We have implemented a general GPU-accelerated nonlinear optimization framework based on Levenberg-Marquardt for sparse estimation problems in CUDA C++, which enables the sharing of data types and functions between real-time applications and optimization tasks.
 Graphite achieves a relatively lower memory footprint by supporting in-place optimization, inferring matrix structures from the optimizable graph, supporting mixed-precision solving, and enabling matrix-free methods.
-Our results show that our solver enables large-scale optimization over complex constraints for desktop and resource-constrained devices.
-Discussion: Currently, Graphite only supports Levenberg-Marquardt and PCG; direct solvers are not yet supported, so the computed parameter step is always inexact, although iterative methods scale better and use less memory.
-It also assumes that all constraints need observations and constraint data, but this is not always the case.
-Still, the overhead of storing these is mitigated by using a small data type (e.g.
-a char).
-Also, the performance could be improved further by extending vertex descriptors to support reduced camera systems, and using vectorized intrinsics and tensor cores.
+Our results show that our framework enables large-scale optimization over complex constraints for desktop and resource-constrained devices.
+In the future, we would like to support more algorithms and solving techniques.